Hi! I'm really having a difficulty answering this question so you might help me with this: solve for the valu of y: by+4=x-2

Answer by Theo(13342)

start with:

by + 4 = x - 2

subtract 4 from both sides of the eqaution to get by = x - 6

divide both sides of the equation by b to get y = (x-6) / b

y is the dependent variable.

x and b are the independent variables.

once you know the values of x and b, you can determine the value of y.